# Robust image representations with counterfactual contrastive learning

### v1

@@ Abstract
Contrastive pretraining can substantially increase model generalisation and downstream performance. However, the quality of the learned representations is highly dependent on the data augmentation strategy applied to generate positive pairs. Positive contrastive pairs should preserve semantic meaning while discarding unwanted variations related to the data acquisition domain. Traditional contrastive pipelines attempt to simulate domain shifts through pre-defined generic image transformations. However, these do not always mimic realistic and relevant domain variations for medical imaging such as scanner differences. To tackle this issue, we herein introduce counterfactual contrastive learning, a novel framework leveraging recent advances in causal image synthesis to create contrastive positive pairs that faithfully capture relevant domain variations. Our method, evaluated across five datasets encompassing both chest radiography and mammography data, for two established contrastive objectives (SimCLR and DINO-v2), outperforms standard contrastive learning in terms of robustness to acquisition shift. Notably, counterfactual contrastive learning achieves superior downstream performance on both in-distribution and on external datasets, especially for images acquired with scanners under-represented in the training set. Further experiments show that the proposed framework extends beyond acquisition shifts, with models trained with counterfactual contrastive learning substantially improving subgroup performance across biological sex.
Introduction
@@
Contrastive learning in medical imaging has emerged as an effective strategy to leverage unlabelled data. This self supervised learning approach has been shown to substantially improve model generalisation across domain shifts as well as reducing the amount of high-quality annotated data needed for training [1, 2, 3, 4]. However, the success of contrastive-based learning is heavily dependent on the positive pair generation pipeline [5]. These positive pairs are typically generated by repeatedly applying pre-defined data augmentations to the original image. As such, changes in the augmentation pipeline have a substantial impact on the quality of the learned representations, ultimately influencing downstream performance and robustness to domain changes [5, 6]. Traditionally, augmentation pipelines developed for natural images have been directly applied to medical imaging, however, this might not be optimal due to the unique challenges and characteristics of how medical scans are acquired. In particular, domain variations are often much larger than subtle class-wise differences. This may lead contrastively learned representations to inadvertently encode these irrelevant acquisition-related variations into the learned representations.
@@
In this work, we aim to improve the robustness of contrastively learned representation against domain shifts, in particular acquisition shift. Acquisition shift is caused by changes in image acquisition protocols (device settings, post-processing software, etc.), and is a major source of dataset shift in the medical imaging domain. We hypothesise that robustness of contrastively learned features against such changes in image characteristics could be improved by simulating domain variations more faithfully in the positive pair creation stage. For this reason, we propose and evaluate ‘counterfactual contrastive learning’, a new contrastive pair generation framework leveraging recent advances in deep generative models for high-quality, realistic counterfactual image generation [7, 8]. Counterfactual generation models allow us to answer ‘what-if’ questions, such as simulating how a mammogram acquired with one device would appear if it would have been acquired with a different device. Specifically, in our proposed counterfactual contrastive framework, we create cross-domain positive pairs, by matching real images with their domain counterfactuals, realistically simulating device changes. Importantly, the proposed approach is agnostic to the choice of the contrastive objective as it only impacts the positive pair creation step. We illustrate the benefits of this approach for two widely-used contrastive learning frameworks: seminal work SimCLR [9] and newly released DINO-V2 [10] objectives. Moreover, to precisely measure the effect of the proposed counterfactual pair generation process, we also compare the proposed approach to a simpler approach where we simply extend the training set with the generated counterfactuals.
@@
Evaluating the proposed counterfactual contrastive learning framework across two medical image modalities, mammography and chest radiographs, on five public datasets and two clinically-relevant classification tasks, we show that our method yields features which are more robust to domain changes. This increased robustness was directly observed in the feature space and, more importantly, by a substantial increase in downstream performance, in particular under limited labels and for domains under-represented at training time. Crucially, these findings hold for both SimCLR and DINO-V2, despite major differences in the training objectives. This paper is an extension of our recent workshop paper [11]. It differs in the following aspects:
@@
  • we previously only considered the SimCLR objective, here we extend our counterfactual contrastive approach to the recently proposed DINO-V2 [10] objective. These new results demonstrate empirically that the proposed approach is versatile and agnostic to the choice of the contrastive objective.
@@
  • while the main focus of this work is robustness to acquisition shift, in this extension we show that the proposed method can be extended beyond this scenario, for example to improve subgroup performance.
@@
Figure 1: We propose a novel counterfactual contrastive pair generation framework for improving robustness of contrastively-learning features to distribution shift. As opposed to solely relying on a pre-defined augmentation generation pipeline $\mathcal{T}$ (as in standard contrastive learning), we propose to combine real images with their domain counterfactuals to create realistic cross-domain positive pairs. Importantly, this proposed approach is independent of the specific contrastive objective employed. The causal image generation model is here represented by the ‘do’ operator. We also compare the proposed method to another approach where we simply extend the training set with the generated counterfactual images without explicit matching with their real counterparts, treating real and counterfactuals as independent training samples. Figure adapted from [11]
Related work
Contrastive learning
@@
Generating pairs of image ‘views’ that share the same underlying meaning (positive pairs) is the core principle of contrastive learning. The contrastive objective then encourages the model to learn similar embeddings for these pairs, while keeping them distinct from the embeddings of unrelated images. A landmark work in this field is SimCLR [9], where positive pairs are generated by applying photometric and geometric transformations to the original image. SimCLR stands out for its simplicity, effectiveness, and widespread adoption, particularly in medical imaging. Azizi et al. [2] for example showed that pre-training models with SimCLR substantially improves downstream performance and robustness to various sources of data shifts. Several methods have proposed refinements to SimCLR, for example BYOL [12], MoCo [13], or most recently DINO(-v2) [14, 10]. DINO uses a self-distillation approach without explicit negative pairs, where a student network learns to match the output of a teacher network, updated via momentum. This method focuses on consistency between the teacher and student outputs, making it less dependent on batch size and augmentations. Importantly positive pairs are generated using more than two views. Representations are instead encouraged to be similar across different global views (larger image crops) and local views (smaller crops). Moreover, DINO relies on vision transformers [15], contrarily to SimCLR which was primarily designed for convolutional networks. Further enhancements were proposed in DINO-v2 [10], with modifications to the loss function to improve stability and performance, encouraging better feature alignment and consistency. Recent work has successfully applied DINO-v2 pre-training to chest radiography, achieving state-of-the-art downstream performance across different tasks [16].
Counterfactual image generation
@@
One goal of counterfactual image generation is to produce ‘counterfactual explanations’, i.e. images depicting the smallest change in the input that would have changed the prediction of a pre-defined classifier [17, 18, 19, 20, 21]. Parallel to this interpretability-centered line of work, others have focused on using generative modelling to synthesise ‘what-if’ images, independently of any external classifier. Seminal work by Pawlowski et al. [22] introduced Deep Structural Causal Models (DSCM) to generate realistic counterfactuals for small resolution images. This framework has been substantially extended by Ribeiro et al. [7], where the authors utilise a hierarchical variational autoencoders (HVAE) for improving image generation, unlocking high-quality high-resolution counterfactual generation, in particular for medical images. While counterfactual image generation models are gaining traction, with some studies showing promising results in imbalanced data augmentation [23, 24] and fairness [25], the potential of these models for enhancing performance on clinically relevant tasks still warrants more exploration.
Combining contrastive learning and counterfactuals
@@
Zhang et al. [26] explored the use of counterfactual text-image pairs in vision-language grounding tasks, defining task-dependent counterfactual pairs for additional supervision signal. Within the context of supervised contrastive graph learning, Yang et al. [27] proposed to generate challenging negative examples using graph counterfactuals. However, the use of image-only counterfactuals for vision contrastive learning remains largely unexplored.
Counterfactual contrastive learning
@@
In this section, we introduce counterfactual contrastive learning, a self-supervised learning paradigm to train image encoders robust to domain variations by leveraging state-of-the-art counterfactual image generation models. Contrastive learning typically uses colour and intensity-based image augmentations to encourage the model to ignore domain-specific image characteristics. However, in medical imaging, the effect of changes in acquisition hardware, device calibration or post-processing software on the final image appearance is highly complex and can not realistically be replicated by those simple handcrafted transformations. To overcome this, in counterfactual contrastive learning we instead use a causal image generation model to simulate realistic domain variations and generate cross-domain contrastive pairs, explicitly encouraging contrastively-learned representations to ignore domain-specific information (such as scanner differences). We illustrate key differences between standard contrastive learning and the proposed counterfactual contrastive approach in Fig. 1.
Counterfactual image generation model
@@
In this work, we use the Deep Structural Causal Model (DSCM) proposed by Ribeiro et al. [7] to generate image counterfactuals. In this model, images are generated using a Hierarchical Variational Autoencoder (HVAE) conditioned on the assumed causal parents of the image. This HVAE is trained by maximising the Evidence Lower Bound of the log-likelihood. We follow Ribeiro et al. and use a conditional prior for the latent representations of the image. Once the HVAE is trained, we can infer the exogenous noise (capturing image-specific identity, not explained by the causal parents) from the observed image and true causal parents. Combining this exogenous noise with the new parents (after intervention) we can then generate a counterfactual image. We invite the reader to consult Ribeiro et al. [7] for a detailed explanation of the counterfactual generation model. A few modifications were made to the original HVAE model from [7] to further increase training stability and image quality. First, instead of directly using the parent variables to condition the HVAE, we add an embedding layer to learn a more flexible parents’ embedding for improved conditioning. Moreover, we used SiLU [28] activation layers instead ReLU, added group normalisation layers and fixed the variance of latent variables to $1e^{-2}$, as we noticed that these changes improved training stability. Finally, our generation model achieved sufficient image quality without relying on any counterfactual finetuning, i.e. our generation model did not need to rely on any pretrained classifiers to reach satisfying levels of effectiveness (which is practical as the use of counterfactual finetuning was recently shown to lead to attribute amplification [29]). Details about the assumed causal graphs and the obtained quality for image generation can be found in the Experiments section.
A simple and effective approach to counterfactual contrastive learning: CF-SimCLR

### v2

@@ Abstract
Contrastive pretraining can substantially increase model generalisation and downstream performance. However, the quality of the learned representations is highly dependent on the data augmentation strategy applied to generate positive pairs. Positive contrastive pairs should preserve semantic meaning while discarding unwanted variations related to the data acquisition domain. Traditional contrastive pipelines attempt to simulate domain shifts through pre-defined generic image transformations. However, these do not always mimic realistic and relevant domain variations for medical imaging, such as scanner differences. To tackle this issue, we herein introduce counterfactual contrastive learning, a novel framework leveraging recent advances in causal image synthesis to create contrastive positive pairs that faithfully capture relevant domain variations. Our method, evaluated across five datasets encompassing both chest radiography and mammography data, for two established contrastive objectives (SimCLR and DINO-v2), outperforms standard contrastive learning in terms of robustness to acquisition shift. Notably, counterfactual contrastive learning achieves superior downstream performance on both in-distribution and external datasets, especially for images acquired with scanners under-represented in the training set. Further experiments show that the proposed framework extends beyond acquisition shifts, with models trained with counterfactual contrastive learning reducing subgroup disparities across biological sex.
1 Introduction
@@
Contrastive learning in medical imaging has emerged as an effective strategy to leverage unlabelled data. This self-supervised learning approach has been shown to substantially improve model generalisation across domain shifts as well as reduce the amount of high-quality annotated data needed for training (Azizi et al., 2021, 2023; Ghesu et al., 2022; Zhou et al., 2023). However, the success of contrastive-based learning is heavily dependent on the positive pair generation pipeline (Tian et al., 2020). These positive pairs are typically generated by repeatedly applying pre-defined data augmentations to the original image. As such, changes in the augmentation pipeline have a substantial impact on the quality of the learned representations, ultimately influencing downstream performance and robustness to domain changes (Tian et al., 2020; Scalbert et al., 2023). Traditionally, augmentation pipelines developed for natural images have been directly applied to medical imaging, however, this might not be optimal due to the unique challenges and characteristics of how medical scans are acquired. In particular, domain variations are often much larger than subtle class-wise differences. This may lead contrastively-learned representations to inadvertently encode these irrelevant acquisition-related variations into the learned representations.
Figure 1: We propose a novel counterfactual contrastive pair generation framework for improving robustness of contrastively-learned features to distribution shift. As opposed to solely relying on a pre-defined augmentation generation pipeline $\mathcal{T}$ (as in standard contrastive learning), we propose to combine real images with their domain counterfactuals to create realistic cross-domain positive pairs. Importantly, this proposed approach is independent of the specific contrastive objective employed. The causal image generation model is represented by the ‘do’ operator. We also compare the proposed method to another approach where we simply extend the training set with the generated counterfactual images without explicit matching with their real counterparts, treating real and counterfactuals as independent training samples. Figure adapted from (Roschewitz et al., 2024).
@@
In this work, we aim to improve the robustness of contrastively-learned representation against domain shifts, in particular acquisition shift. Acquisition shift is caused by changes in image acquisition protocols (device settings, post-processing software, etc.), and is a major source of dataset shift in the medical imaging domain. We hypothesise that the robustness of contrastively-learned features against such changes in image characteristics could be improved by simulating domain variations more faithfully in the positive pair creation stage. For this reason, we propose and evaluate ‘counterfactual contrastive learning’, a new contrastive pair generation framework leveraging recent advances in deep generative models for high-quality, realistic counterfactual image generation (Ribeiro et al., 2023; Fontanella et al., 2023). Counterfactual generation models allow us to answer ‘what-if’ questions, such as simulating how a mammogram acquired with one device would appear if it had been acquired with a different device. Specifically, in our proposed counterfactual contrastive framework, we create cross-domain positive pairs, by matching real images with their domain counterfactuals, realistically simulating device changes. Importantly, the proposed approach is agnostic to the choice of the contrastive objective as it only impacts the positive pair creation step. We illustrate the benefits of this approach for two widely-used contrastive learning frameworks: seminal work SimCLR (Chen et al., 2020) and newly released DINO-V2 (Oquab et al., 2023) objectives. Moreover, to precisely measure the effect of the proposed counterfactual pair generation process, we also compare the proposed approach to a simpler approach where we simply extend the training set with the generated counterfactuals.
@@
Evaluating the proposed counterfactual contrastive learning framework across two medical image modalities, mammography and chest radiographs, on five public datasets and two clinically-relevant classification tasks, we show that our method yields features which are more robust to domain changes. This increased robustness was directly observed in the feature space and, more importantly, by a substantial increase in downstream performance, in particular under limited labels and for domains under-represented at training time. Crucially, these findings hold for both SimCLR and DINO-V2, despite major differences in the training objectives. This paper is an extension of our recent MICCAI workshop paper (Roschewitz et al., 2024). It differs in the following aspects:
@@
  • we previously only considered the SimCLR objective, here we extend our counterfactual contrastive approach to the recently proposed DINO-V2 (Oquab et al., 2023) objective. These new results demonstrate empirically that the proposed approach is versatile and agnostic to the choice of the contrastive objective.
@@
  • While the main focus of this work is robustness to acquisition shift, in this extension, we show that the proposed method can be extended beyond this scenario, for example, to improve subgroup performance.
@@
2 Related work
2.1 Contrastive learning
@@
Generating pairs of image ‘views’ that share the same underlying meaning (positive pairs) is the core principle of contrastive learning. The contrastive objective then encourages the model to learn similar embeddings for these pairs, while keeping them distinct from the embeddings of unrelated images. A landmark work in this field is SimCLR (Chen et al., 2020), where positive pairs are generated by applying photometric and geometric transformations to the original image. SimCLR stands out for its simplicity, effectiveness, and widespread adoption, particularly in medical imaging. Azizi et al. (2023) for example showed that pre-training models with SimCLR substantially improves downstream performance and robustness to various sources of data shifts. Several methods have proposed refinements to SimCLR, for example BYOL (Grill et al., 2020), MoCo (He et al., 2020), or most recently DINO(-v2) (Caron et al., 2021; Oquab et al., 2023). DINO uses a self-distillation approach without explicit negative pairs, where a student network learns to match the output of a teacher network, updated via momentum. This method focuses on consistency between the teacher and student outputs, making it less dependent on batch size and augmentations. Importantly positive pairs are generated using more than two views. Representations are instead encouraged to be similar across different global views (larger image crops) and local views (smaller crops). Moreover, DINO relies on vision transformers (Dosovitskiy et al., 2020), contrarily to SimCLR which was primarily designed for convolutional networks. Further enhancements were proposed in DINO-v2 (Oquab et al., 2023), with modifications to the loss function to improve stability and performance, encouraging better feature alignment and consistency. Recent work has successfully applied DINO-v2 pre-training to chest radiography, achieving state-of-the-art downstream performance across different tasks (Moutakanni et al., 2024).
Contrastive learning is a popular paradigm in self-supervised learning for medical imaging, in particular SimCLR which was found to be the most used contrastive objective in a recent review by Huang et al. (2023). While most works use standard positive pair creation (designed for natural images) (Azizi et al., 2021, 2023; Moutakanni et al., 2024), some have derived medical imaging-specific positive pair creation methods such as using neighbouring slices in CT images (Dong et al., 2021), neighbouring patches in histopathology (Li et al., 2021), leveraging multiple views in mammography (Ghosh et al., 2024), or using patient metadata to find similar images (Vu et al., 2021; Dufumier et al., 2021). Another approach consists of leveraging multi-modal patient information to form positive pairs, e.g. reports and chest X-rays (Zhang et al., 2022; Ghosh et al., 2024) or combining imaging with genetic information (Taleb et al., 2021). However, such paired views or multi-modal pairs may not always be available, further motivating our investigation of the use of synthetic pairs for medical image contrastive learning.
Previously, von Kügelgen et al. (2021) have theoretically demonstrated that contrastive learning provably disentangles content from style, provided that augmentations capture realistic style changes, proposing to interpret data augmentation in contrastive settings as ‘counterfactuals under soft style intervention’ and providing a strong theoretical motivation to our work.
2.2 Counterfactual image generation
@@
One goal of counterfactual image generation is to produce ‘counterfactual explanations’, i.e. images depicting the smallest change in the input that would have changed the prediction of a pre-defined classifier (Augustin et al., 2022; Sanchez et al., 2022; Atad et al., 2022; Matsui et al., 2022; Sun et al., 2024). Parallel to this interpretability-centered line of work, others have focused on using generative modelling to synthesise ‘what-if’ images, independently of any external classifier. Seminal work by Pawlowski et al. (2020) introduced Deep Structural Causal Models (DSCM) to generate realistic counterfactuals for small-resolution images. This framework has been substantially extended by Ribeiro et al. (2023), where the authors utilise a hierarchical variational autoencoder (HVAE) for improving image generation, unlocking high-quality high-resolution counterfactual generation, in particular for medical images. While counterfactual image generation models are gaining traction, with some studies showing promising results in imbalanced data augmentation (Xia et al., 2022; Garrucho et al., 2023) and fairness (Dash et al., 2022), the potential of these models for enhancing performance on clinically relevant tasks still warrants more exploration.
2.3 Combining contrastive learning and counterfactuals
@@
Zhang et al. (2020) explored the use of counterfactual text-image pairs in vision-language grounding tasks, defining task-dependent counterfactual pairs for additional supervision signal. Within the context of supervised contrastive graph learning, Yang et al. (2023) proposed to generate challenging negative examples using graph counterfactuals. However, the use of image-only counterfactuals for vision contrastive learning remains largely unexplored.
3 Counterfactual contrastive learning
@@
In this section, we introduce counterfactual contrastive learning, a self supervised learning paradigm to train image encoders robust to domain variations by leveraging state-of-the-art counterfactual image generation models. Contrastive learning typically uses colour and intensity-based image augmentations to encourage the model to ignore domain-specific image characteristics. However, in medical imaging, the effect of changes in acquisition hardware, device calibration or post-processing software on the final image appearance is highly complex and can not realistically be replicated by those simple handcrafted transformations. To overcome this, in counterfactual contrastive learning, we instead use a causal image generation model to simulate realistic domain variations and generate cross-domain contrastive pairs, explicitly encouraging contrastively-learned representations to ignore domain-specific information (such as scanner differences). We illustrate key differences between standard contrastive learning and the proposed counterfactual contrastive approach in Fig. 1.
3.1 Counterfactual image generation model
Formally, a Structural Causal Model (Pearl, 2009) (SCM) is defined by a triple $\mathcal{M}=\langle\mathbf{X},\mathbf{U},\mathbf{F}\rangle$, where $\mathbf{X}=\{X_{i}\}_{i=1}^{n}$ represents the set of endogenous (observed) variables, $\mathbf{U}=\{U_{i}\}_{i=1}^{n}$ a set of exogenous (unobserved) variables, and $\mathbf{F}=\{f_{i}\}_{i=1}^{n}$ a set of functions, aka causal mechanisms, such that $X_{k}:=f_{k}(\mathbf{pa}_{k},U_{k})$, where $\mathbf{pa}_{k}\subseteq\mathbf{X}\setminus X_{k}$ are called parents (i.e. direct causes) of each $X_{k}$. The variables in $\mathbf{X}$ are called endogenous since they are caused by the variables in the model, whereas variables in $\mathbf{U}$ are exogenous as they are caused by factors which are external to the model. The process of generating counterfactuals can then be divided into three steps:
  • Abduction: infer the posterior exogenous noise distribution $P_{\mathbf{U}\mid\mathbf{X}}$ given observed evidence $\mathbf{X}$;
  • Intervention: perform an intervention by modifying one or more of the endogenous variables, e.g. $do(X_{k}:=x)$, to obtain a modified SCM $\mathcal{M}_{x}$;
  • Prediction: use $\mathcal{M}_{x}$ and $P_{\mathbf{U}\mid\mathbf{X}}$ to compute a counterfactual.
@@ 3.1 Counterfactual image generation model
Here, we use the Deep Structural Causal Model (DSCM) proposed by Ribeiro et al. (2023) to generate image counterfactuals. In what follows, we detail the counterfactual image generation process using this model. For simplicity, we here assume no further causal relationships between the parents of the image (i.e. all image parents are independent of each other). Formally, let $\mathbf{x}$ be the image and $\mathbf{pa_{x}}$ the parents of $\mathbf{x}$. In Ribeiro et al. (2023), the mechanism $\mathbf{x}\coloneqq f_{\mathbf{x}}(\mathbf{pa}_{\mathbf{x}},\mathbf{u}_{\mathbf{x}})$ is modelled using an HVAE. Specifically, the exogenous noise is decomposed into two parts $p(\mathbf{u}_{\mathbf{x}})=p_{\theta}(\mathbf{z})p(\boldsymbol{\epsilon})$, where the first is inferred using the HVAE’s encoder $q_{\phi}(\mathbf{z}\mid\mathbf{x},\mathbf{pa}_{\mathbf{x}})$, and the second by inverting the decoder $g_{\theta}$’s sampling mechanism $\boldsymbol{\epsilon}=(\mathbf{x}-\boldsymbol{\mu}(\mathbf{z},\mathbf{pa}_{\mathbf{x}}))/\boldsymbol{\sigma}(\mathbf{z},\mathbf{pa}_{\mathbf{x}})$, where $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$ here are per pixel mean/std predictions by the decoder. To generate a counterfactual image $\widetilde{\mathbf{x}}$, we simply compute $\widetilde{\mathbf{x}}=\boldsymbol{\mu}(\mathbf{z},\widetilde{\mathbf{pa}}_{\mathbf{x}})+\boldsymbol{\sigma}(\mathbf{z},\widetilde{\mathbf{pa}}_{\mathbf{x}})\odot\boldsymbol{\epsilon}$ holding the exogenous noise fixed, where $\widetilde{\mathbf{pa}}_{\mathbf{x}}$ are the counterfactual parents obtained after modifying one or more of $\mathbf{pa}_{\mathbf{x}}$’s values. Note that in this work, we use the conditional prior proposed by Ribeiro et al. (2023) in the HVAE, i.e. $p_{\theta}(\mathbf{z}_{1:L}\mid\mathbf{pa}_{\mathbf{x}})=p_{\theta}(\mathbf{z}_{L}\mid\mathbf{pa}_{\mathbf{x}})\prod_{i=1}^{L-1}p_{\theta}(\mathbf{z}_{i}\mid\mathbf{z}_{>i},\mathbf{pa}_{\mathbf{x}}).$ The HVAE is then trained by maximising the Evidence Lower Bound (ELBO) of the log-likelihood on the observed dataset. In this work, minor modifications were made to the original HVAE model from Ribeiro et al. (2023) to further increase training stability and image quality. First, instead of directly using the parent variables to condition the HVAE, we add an embedding layer to learn a more flexible parents’ embedding for improved conditioning. Moreover, we used SiLU (Hendrycks and Gimpel, 2023) activation layers instead of ReLU, added group normalisation layers, and fixed the decoder’s variance to 1e-2, as we noticed that these changes improved training stability.
Note that, in their initial study, Ribeiro et al. (2023) noticed that solely relying on likelihood training for the HVAE may sometimes lead to ‘ignored counterfactual conditioning’ i.e. $\mathbf{\widetilde{x}}$ does not obey $\mathbf{\widetilde{pa}_{x}}$. To mitigate this issue, the authors introduced counterfactual finetuning. This step is optional and involves further finetuning the HVAE weights after the likelihood training in order to improve effectiveness of the model (i.e. how well the generated counterfactual obeys the counterfactual parents). Concretely, this method leverages a pre-trained classifier $q_{\psi}(\mathbf{\widetilde{pa}_{x}}\mid\widetilde{\mathbf{x}})$, and optimises the pre-trained HVAE weights $\{\theta,\phi\}$ to maximise $\log q_{\psi}(\mathbf{\widetilde{pa}_{x}}\mid\widetilde{\mathbf{x}})$ with $\psi$ fixed. However, Xia et al. (2024) showed that this finetuning step may overly emphasise non-intervened attributes, a phenomenon known as ‘attribute amplification’, and it can be mitigated by using soft rather than hard labels during finetuning. In this study, we found that our generation model achieved sufficient image quality without relying on any counterfactual finetuning. Hence, unless otherwise stated, the counterfactual image generation used in the following does not rely on this step.
3.2 A simple and effective approach to counterfactual contrastive learning: CF-SimCLR